OpenStudy (anonymous):

please help and explain not just answer y^-2*y^6 simplify positive exponent only

Nnesha (nnesha):

when we multiply same bases we should `add` their exponents \[\huge\rm x^m \times x^n=x^{m+n}\]

Nnesha (nnesha):

so \[\rm y^{-2} \times y^6= ??\]

OpenStudy (anonymous):

y^4 is what i got and it was correct

Nnesha (nnesha):

yes that's right -2+(6) =4

OpenStudy (anonymous):

ty

Nnesha (nnesha):

np :=)